The schematic shows a steam power plant in which the turbine develops 1050 kW. The heat supplied to the steam in the boiler is 5040 kJ/kg, the heat rejected by the steam to the cooling water in the condenser is 3780 kJ/kg and the feed pump work required to pump the condensate back into the boiler is 6 kW.

Calculate:

a. the steam flow rate in kg/s.

b. The amount of heat transferred for 10 kg of steam. Follow your book’s sign convention.


1
Expert's answer
2020-11-16T07:48:40-0500

a. "Q_{Net} = W_{Net}"

"m_{sfr} \times (5040 – 3780) = (1050 – 6)"

"m_{sfr} = \frac{1044}{1260} = 0.82 \;kg/s"

b. "Q_{Net} = m \times 5040 = 10 \times 5400 = 54000 \;kJ"